Use a suitable identity to get each of the following products.
(3a-1/2)(3a- 1/2)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(3a-1/2)(3a- 1/2)
= (3a-1/2)(3a-1/2)= 9a^2-3a+(1/4)
Using (a-b)^2= a^2+ b^2- 2ab

Thanks